ae (1.2). ee (1.2.34). a (1.2.34,5,6). aE1.2.3,5,67 a (2.3.7,8,15, 17.24.26 (2,3.524.73,221,225,627), p= 1237 -3 p29


Quadratic Reciprocity ) Compute () (We know many different ways to complete this computation: beute foece, Euler's Criterion Gauss Lemma, Quadratic Recipeocity. Be sre to nderstand how to apply all of these methods)
ae (1.2). ee (1.2.34). a (1.2.34,5,6). aE1.2.3,5,67 a (2.3.7,8,15, 17.24.26 (2,3.524.73,221,225,627), p= 1237 -3 p29
Compucte (we no mang different wos to complete this Computafion bute Foce Buler's Crterion Gauss' lenma Cuadratic Reciprocity. Be sure ts undavstond has t apply al these muthods) -a E L21 a 6 2,,549] p aet 2.3.78.s,A, 14,14) ps aG ibs 24, 41
Answer #1

We have to compute a . Basic rules we will follow are as follows:

1. \left(\dfrac{a}{p}\right)=\left(\dfrac{b}{p}\right) if а — 6(d p) тос

2.\left(\dfrac{p}{q}\right) = \left(\dfrac{q}{p}\right) if p 1(mod 4) or q 1(mod 4) .​

Otherwise, if p 3(mod 8) and q 3(mod 8) , we have P P .

3. p-1)/2 (mod p) a P

4. 2 = (-1)(p°-1)/8 (mod p) t1(mod 8) 1(if p and = 1(if p+3(mod 8))

5.\left(\dfrac{x^2}{p}\right) = 1 (if~p\nmid x) ,=0 ~otherwise

6. \left(\dfrac{ab}{p}\right) = \left(\dfrac{a}{p}\right)\left(\dfrac{b}{p}\right)

Using the above things, computations can be done as follows:

\left(\dfrac{1}{3}\right) = 1^{(3-1)/2}(mod~3)=1

\left(\dfrac{2}{3}\right) = -1 since 3 3(mod 8)

\left(\dfrac{1}{5}\right) = 1^{(5-1)/2}(mod~3)=1

\left(\dfrac{2}{5}\right) = -1 since 5-3(mod 8)

\left(\dfrac{3}{5}\right) = 3^{(5-1)/2}(mod~ 5)=3^2(mod~5)=9(mod~5) = -1

\left(\dfrac{4}{5}\right)=\left(\dfrac{2^2}{5}\right)=1 since 5\nmid2

\left(\dfrac{1}{7}\right) = 1

\left(\dfrac{2}{7}\right) = 1 (\because 7\equiv-1(mod~8))

\left(\dfrac{3}{7}\right) = -\left(\dfrac{7}{3}\right) (\because 3 \equiv3(mod~4) ~and ~7\equiv3(mod~4))=-\left(\dfrac{1}{3}\right) =-(1)=-1

\left(\dfrac{5}{7}\right) = \left(\dfrac{7}{5}\right) (\because 5\equiv1(mod~4))\\~~~~~~~~~~~~=\left(\dfrac{2}{5}\right)(\because 7\equiv2(mod~5))\\~~~~~~~~~~~~=-1(\because5\equiv-3(mod~8))

\left(\dfrac{6}{7}\right) = \left(\dfrac{2}{7}\right)\left(\dfrac{3}{7}\right) = (1)(-1)=-1

\left(\dfrac{1}{11}\right) = 1^{(11-1)2}=1

\left(\dfrac{2}{11}\right) = -1 ~(\because11\equiv3(mod~8))

\left(\dfrac{3}{11}\right) = - \left(\dfrac{11}{3}\right)(\because 11\equiv3(mod~4)~and~3\equiv3(mod~4))\\~~~~~~~~~~~~~~=-\left(\dfrac{2}{3}\right)(\because11\equiv2(mod~3))\\~~~~~~~~~~~~~~=-(-1) (\because 3\equiv3(mod~8))\\~~~~~~~~~~~~~~=1

\left(\dfrac{5}{11}\right) = \left(\dfrac{11}{5}\right)(\because 5\equiv1(mod~4))\\~~~~~~~~~~~~~~=\left(\dfrac{1}{5}\right)(\because11\equiv1(mod~5))\\~~~~~~~~~~~~~~=1

\left(\dfrac{6}{11}\right) = \left(\dfrac{2}{11}\right)\left(\dfrac{3}{11}\right)\\~~~~~~~~~~~~~~=(-1)(1)\\~~~~~~~~~~~~~~=-1

7 (. 11 3(mod 4) and 7 3(mod 4) 7 11 (114(mod 11)) 22 7 -(1 72)

Four of the given questions are answered. The remaining can be computed similarly.

I will demonstrate the last one since the numbers seem very large.

\left(\dfrac{627}{1237}\right) = \left(\dfrac{1237}{627}\right) (\because 1237\equiv1(mod~4))\\~~~~~~~~~~~~~~~~~=\left(\dfrac{610}{627}\right)(\because 1237\equiv610(mod~627))\\~~~~~~~~~~~~~~~~~= \left(\dfrac{2}{627}\right)\left(\dfrac{5}{627}\right)\left(\dfrac{61}{627}\right)\\~~~~~~~~~~~~~~~~~=(-1)\left(\dfrac{627}{5}\right)\left(\dfrac{627}{61}\right)(\because 627\equiv3(mod~8), 5\equiv1(mod~4),61\equiv1(mod~4))\\~~~~~~~~~~~~~~~~~~=(-1)\left(\dfrac{2}{5}\right)\left(\dfrac{17}{61}\right)(\because 627\equiv2(mod~5),627\equiv17(mod~61))\\~~~~~~~~~~~~~~~~~~=(-1)(-1)\left(\dfrac{61}{17}\right)(\because5\equiv3(mod~8),17\equiv1(mod~4))\\~~~~~~~~~~~~~~~~~~=\left(\dfrac{10}{17}\right)(\because 61\equiv10(mod~17))\\~~~~~~~~~~~~~~~~~~=\left(\dfrac{2}{17}\right)\left(\dfrac{5}{17}\right)=(1)\left(\dfrac{17}{5}\right)(\because17\equiv1(mod~8) ,17\equiv1(mod~4))\\£~~~~~~~~~~~~~~~~~~=\left(\dfrac{2}{5}\right)(\because17\equiv2(mod~5))\\~~~~~~~~~~~~~~~~~~=-1(\because 5\equiv-3(mod~8)

Just by following the same technique, remaining can be calculated.